OpenStudy (anonymous):

how to find particular solution of differential equation. y=23/x; y=16 when x=1

OpenStudy (anonymous):

Do you mean \(y'=\dfrac{23}{x}\) ?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

You're given a separable equation. Split the differentials: \[\frac{dy}{dx}=\frac{23}{x}~~\iff~~dy=\frac{23}{x}~dx\] Then integrate both sides.

OpenStudy (anonymous):

can you write it out for me please?

OpenStudy (anonymous):

\[\int dy=\int\frac{23}{x}~dx\]

OpenStudy (anonymous):

nxt steps?

OpenStudy (jhannybean):

\[\frac{d}{dx} \ln(x) = \frac{1}{x} \iff \int \frac{1}{x} dx =~?\]

OpenStudy (anonymous):

lnx+c whats next?